OpenStudy (anonymous):

Calculus- integration 4/(x^6-8 x^3) find partial fraction decomp

OpenStudy (anonymous):

I need help finding the decomposition. Please

OpenStudy (freckles):

hello

OpenStudy (freckles):

can you factor the bottom any?

OpenStudy (anonymous):

hi freckles..again

OpenStudy (anonymous):

the bottom is what has me confused im thinking pull out \[x ^{3}\]

OpenStudy (anonymous):

if i did that then it would be \[x^3 (x^2 - 8)\] is that correct?

OpenStudy (freckles):

almost

OpenStudy (freckles):

x^3*x^3=x^(3+3)=x^6 correct?

OpenStudy (freckles):

\[\frac{4}{x^6-8x^3}=\frac{4}{x^3(x^3-8)}=\frac{4}{x^3(x-2)(x^2+2x+4)}\]

OpenStudy (freckles):

do you know how I factor x^3-8 ?

OpenStudy (anonymous):

2^3 = 8 so you pull out a factor of x-2

OpenStudy (freckles):

well it was just a difference of cubes

OpenStudy (anonymous):

after the x-2 did you use long division again here?

OpenStudy (anonymous):

oh

OpenStudy (freckles):

you could if you didn't know the difference of cubes formula

OpenStudy (anonymous):

im gonna have to look that up

OpenStudy (freckles):

\[\text{ Difference of cubes } \\ u^3-a^3=(u-a)(u^2+ua+a^2) \\ \text{ Sum of cubes } \\ u^3+a^3=(u+a)(u^2-ua+a^2)\]

OpenStudy (anonymous):

the numerator..... would that be A + B + Cx+D ?

OpenStudy (freckles):

Well the thing is we have a linear repeated three times And we have another linear after that (with no repetition ) and a quadratic

OpenStudy (freckles):

for a linear the numerator should be a constant for a quadratic the numerator should be a linear

OpenStudy (anonymous):

so the x^3 would be Ax+B

OpenStudy (freckles):

\[\frac{4}{x^3(x-2)(x^2+2x+4)} \\ =\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x^3}+\frac{D}{x-2}+\frac{Ex+F}{x^2+2x+4}\]

OpenStudy (anonymous):

oh wow

OpenStudy (anonymous):

because for the x^3.. we have to show the other factors?

OpenStudy (freckles):

well don't look at x^2 and x^3 as a quadratic and cubic respectively look at them as linears raise to a power

OpenStudy (anonymous):

why doesnt the X^2 and x^3 have Ax+B in the numerator is the degree is higher than 1?

OpenStudy (anonymous):

i thought anything over x^1 would need the Ax+B in the numerator. This is what really confuses me

OpenStudy (freckles):

it is because x is linear

OpenStudy (freckles):

the thing inside the power is linear

OpenStudy (anonymous):

oh.. ok!

OpenStudy (anonymous):

so thhe Ex+F..... that was given the x and 2 coefficients because it can not factor?

OpenStudy (freckles):

linears go over irreducible quadratics

OpenStudy (anonymous):

my problem occurs with the numerator and I get confused to whether or not it will need the Ax+B in it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks again

OpenStudy (freckles):

\[\text{ pretend we have } \frac{x+1}{x^3}\] like we aren't go to look at the bottom as a cubic we are going to look at as a linear cubed

OpenStudy (anonymous):

ok..

OpenStudy (freckles):

\[\frac{x+1}{x^3}=\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x^3}\]

OpenStudy (anonymous):

ok

OpenStudy (freckles):

if we had \[\frac{x+1}{(x^2+1)^3}=\frac{Ax+B}{x^2+1}+\frac{Cx+D}{(x^2+1)^2}+\frac{Ex+F}{(x^2+1)^3}\]

OpenStudy (freckles):

that is because we had a irreducible quadratic cubed

OpenStudy (anonymous):

ooohhh I see

OpenStudy (anonymous):

i see. and x by itself, no matter the degree...is considered linear

OpenStudy (freckles):

\[\frac{x+1}{x^3(x^2+1)^3}=\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x^3}+\frac{Dx+E}{x^2+1}+\frac{Fx+G}{(x^2+1)^2}+\frac{Hx+J}{(x^2+1)^3}\]

OpenStudy (anonymous):

oh my gosh you are soo helpful! I love when your on here

OpenStudy (freckles):

yes

OpenStudy (anonymous):

thanks a bunch!

OpenStudy (freckles):

one more example though just for fun so you can determine a linear from a quadratic or whatever

OpenStudy (anonymous):

k

OpenStudy (freckles):

\[\frac{x}{(x-2)^2(x-1)(x^2+1)^2}\] how would you set this one up

OpenStudy (anonymous):

i am slow at the equation thing, so gimme a min to type it in lol

OpenStudy (freckles):

it's cool

OpenStudy (anonymous):

\[A/ (x-2) + B/ (x-2)^+ C/(x-1)+ Dx+E/(x^2+1) + Fx+G/ (x^2 +1)^2\]

OpenStudy (anonymous):

i get confused when im typing

OpenStudy (anonymous):

this is way wrong huh?

OpenStudy (freckles):

I think you just said this and if you did I'm totally totally proud of you \[\frac{A}{x-2}+\frac{B}{(x-2)^2}+\frac{C}{x-1}+\frac{Dx+E}{x^2+1}+\frac{Fx+G}{(x^2+1)^2}\] you missed the power on the second fraction but I think you meant to put it :)

OpenStudy (anonymous):

that plus sign is supposed to be a 2 :(

OpenStudy (anonymous):

so i did it right?

OpenStudy (freckles):

yep

OpenStudy (anonymous):

awesome! I thank you kindly. Can i Private Message you in the future? You are so helpful to me

OpenStudy (freckles):

sure

OpenStudy (anonymous):

great!